# Approximating the coefficients of the Bessel functions

### v2

@@ 10.2. Upper bounds of the magnitudes of Bessel functions
In preparation for studying the uniform convergence of Bessel functions, we prove some upper bounds on the functions assuming that $\theta\in\theta(\mathcal{R})$ is nonnegative so that we can apply Theorem˜1.2.
@@
First, observe that from Theorem˜1.2,
@@
(A): We use Theorem˜1.2. Note that for $z\in\mathbb{C}^{N}$, we let $\text{Re}(z)$ denote $(\text{Re}(z_{i}))_{i\in[N]}$ and for $r\in\mathbb{R}^{N}$, we let $|r|$ denote $(|r_{i}|)_{i\in[N]}$. First, observe that
@@
Therefore, after referring to Theorem˜1.2, we can deduce that
@@
    Assume that $\theta_{0},\theta_{1}\in\mathbb{C}$ such that $\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))$ is invertible. The value of $J_{a}^{BC^{N}(\theta_{0},\theta_{1})}(z,0\ldots,0)$ is $1+\sum_{k=1}^{\infty}\frac{c_{k}(a^{2})}{4^{k}(\theta_{1}+(N-1)\theta_{0}+\frac{1}{2})_{k}(\theta_{0}N)_{k}}z^{2k}$.
@@
$J_{a}^{A^{N-1}(\theta)}(z,0\ldots,0)=1+\sum_{k=1}^{\infty}r_{k}(a)z^{k}.$
@@ 10.4. The uniform convergence of Bessel functions for Vershik-Kerov
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$. Assume that $\theta\in\mathbb{C}$. Consider the following statements.
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$. Assume that $\theta_{0}\in\mathbb{C}$ and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$. Consider the following statements.
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$. Assume that $\theta\in\mathbb{C}$. Consider the following statements.
@@
(A): We include the proof of this result from [rank_infinity] for completeness. Using Theorem˜1.2 and Hölder’s inequality gives that
@@
It is then is clear that $\zeta\triangleq\{J_{a(N)}^{A^{N-1}(\theta)}(x_{1},\ldots,x_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{rM}\,\forall i\in[r]\}$. Following this, we can use Corollary˜9.2 to deduce that the limit of the coefficient of $r(x)$ in $\zeta$ equals its coefficient in $\exp\left(\sum_{k\geq 1}\frac{m_{k}p_{(k)}\theta^{-k+1}}{k}\right)$. We can then apply part (B) of Lemma˜10.8 to conclude the proof.
@@ 10.5. Bessel generating functions for exponentially decaying probability measures
Suppose $\mu$ is a probability distribution over $\mathbb{C}^{N}$. Then, the Bessel generating function associated to $\mu$ and $\mathcal{R}(\theta)$ is
@@
Lemma 10.16 ((A) is [limits_prob_measures]*Lemma 1.4 and (C) is [matrix]*Proposition 2.11).
@@
    There exist unique constants $c_{\lambda}(\mu)\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$ such that
    $G_{\mu}^{\mathcal{R}(\theta)}(x)=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(\mu)p_{\lambda}\right)$
@@
    $1]\mathcal{D}(\mathcal{R}(\theta))(p_{\lambda})G_{\mu}^{\mathcal{R}(\theta)}(x)=\operatorname{\mathbb{E}}_{a\sim\mu}[p_{\lambda}(a)].$
@@
For (B), we note that $G_{\mu}^{\mathcal{R}(\theta)}(x)$ is holomorphic and $G_{\mu}^{\mathcal{R}(\theta)}(0)=1$. ∎
@@
The papers [matrix, limits_prob_measures] only consider the $A^{N}$ root system. However, (A) and (C) are generalizable to any finite root system after using the method discussed in [limits_prob_measures] that involves applying the results of [dunklbound].
@@
Define $c_{\lambda}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$ as in part (B) of Lemma˜10.16. Then, the following are equivalent.
@@
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.

### v3

@@ 10.2. Upper bounds of the magnitudes of Bessel functions
In preparation for studying the uniform convergence of Bessel functions, we prove some upper bounds on the functions assuming that $\theta\in\theta(\mathcal{R})$ is nonnegative so that we can apply Theorem˜1.4.
@@
First, observe that from Theorem˜1.4,
@@
(A): We use Theorem˜1.4. Note that for $z\in\mathbb{C}^{N}$, we let $\text{Re}(z)$ denote $(\text{Re}(z_{i}))_{i\in[N]}$ and for $r\in\mathbb{R}^{N}$, we let $|r|$ denote $(|r_{i}|)_{i\in[N]}$. First, observe that
@@
Therefore, after referring to Theorem˜1.4, we can deduce that
@@
    Assume that $\theta_{0},\theta_{1}\in\mathbb{C}$ such that $\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))$ is invertible. The value of $J_{a}^{BC^{N}(\theta_{0},\theta_{1})}(z,0,\ldots,0)$ is $1+\sum_{k=1}^{\infty}\frac{c_{k}(a^{2})}{4^{k}(\theta_{1}+(N-1)\theta_{0}+\frac{1}{2})_{k}(\theta_{0}N)_{k}}z^{2k}$.
@@
$J_{a}^{A^{N-1}(\theta)}(z,0,\ldots,0)=1+\sum_{k=1}^{\infty}r_{k}(a)z^{k}.$
@@ 10.4. The uniform convergence of Bessel functions for Vershik-Kerov
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$. Assume that $\theta\in\mathbb{C}$ is fixed. Consider the following statements.
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$. Assume that $\theta_{0}\in\mathbb{C}$ is fixed and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$. Consider the following statements.
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$ for all $N\geq 1$. Assume that $\theta\in\mathbb{C}$ is fixed. Consider the following statements.
@@
(A): We include the proof of this result from [rank_infinity] for completeness. Using Theorem˜1.4 and Hölder’s inequality gives that
@@
It is then is clear that $\zeta\triangleq\{J_{a(N)}^{A^{N-1}(\theta)}(x_{1},\ldots,x_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\{z\in\mathbb{C}^{r}:|\text{Re}(z_{i})|<\frac{\theta}{rM}\,\forall i\in[r]\}$. Following this, we can use Corollary˜9.2 to deduce that the limit of the coefficient of $r(x)$ in $\zeta$ equals its coefficient in $\exp\left(\sum_{k\geq 1}\frac{m_{k}p_{(k)}\theta^{-k+1}}{k}\right)$; the application of Corollary˜9.2 is the difference between this proof and that of [rank_infinity]. We can then apply part (B) of Lemma˜10.8 to conclude the proof.
@@ 10.5. Bessel generating functions for exponentially decaying probability measures
Suppose $\mu$ is a Borel probability distribution over $\mathbb{C}^{N}$. Then, the Bessel generating function associated to $\mu$ and $\mathcal{R}(\theta)$ is
@@
Lemma 10.16 ((A) is [limits_prob_measures]*Lemma 1.4 and (E) is [matrix]*Proposition 2.11).
@@
    There exist unique constants $c_{\lambda}^{\mathcal{R}(\theta)}(\mu)\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$ such that
    $G_{\mu}^{\mathcal{R}(\theta)}(x)=\exp\left(\sum_{\lambda\in\Gamma_{N}}c_{\lambda}^{\mathcal{R}(\theta)}(\mu)p_{\lambda}\right)$
@@
    There exist unique constants $m_{\epsilon}^{\mathcal{R}(\theta)}(\mu)\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{N}$ such that
    $G_{\mu}^{\mathcal{R}(\theta)}(x)=\exp\left(\sum_{\epsilon\in\mathcal{P}_{N}}m_{\epsilon}^{\mathcal{R}(\theta)}(\mu)M_{\epsilon}\right)$
    in a neighborhood of the origin.
  4. (D)
    For all $\lambda\in\Gamma$, the function $\mathcal{D}(\mathcal{R}(\theta))(p_{\lambda})G_{\mu}^{\mathcal{R}(\theta)}(x)$ is holomorphic and equals
    $\int_{\mathbb{C}^{N}}p_{\lambda}(a)J_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)$
    over the open ball of radius $R$ centered at the origin.
  5. (E)
@@
    $\mathcal{D}(\mathcal{R}(\theta))(p_{\lambda})G_{\mu}^{\mathcal{R}(\theta)}(0)=\operatorname{\mathbb{E}}_{a\sim\mu}[p_{\lambda}(a)].$
@@
For (B) and (C), we note that $G_{\mu}^{\mathcal{R}(\theta)}(x)$ is holomorphic in a neighborhood of the origin and $G_{\mu}^{\mathcal{R}(\theta)}(0)=1$.
(D): We include a detailed proof of this statement for completeness. We prove this result for the nonsymmetric eigenfunction $E_{a}^{\mathcal{R}(\theta)}(x)$. It is straightforward to deduce the result for the Bessel function afterwards using the expression relating the two functions in Theorem˜2.8. Note that some of the ideas of this proof have previously appeared in [limits_prob_measures]*Remark 4.6.
Suppose $x\in\overline{B}_{R}(0)$, which is the closed ball of radius $R$ centered at the origin. By part (e) of Lemma˜2.6, the integral
$\int_{\mathbb{C}^{N}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)$
converges because $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\left\lVert a\right\rVert_{2}\left\lVert x\right\rVert_{2})$ and $\mu$ exponentially decays at rate $R$. Furthermore, we can show that the integral is holomorphic over $\overline{B}_{R}(0)$ by following [matrix]*Proof of Lemma 2.9. Note that we have just proved the nonsymmetric version of (A), which can be used to prove (A).
The goal is to prove using induction that for any sequence $\{i_{j}\}_{1\leq j\leq m}$ of elements of $[N]$, we have that for $x\in B_{R}(0)$,
$\prod_{j=1}^{m}\mathcal{D}_{i_{j}}\int_{\mathbb{C}^{N}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)=\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)$
and both sides of the equation converge to a holomorphic function over $B_{R}(0)$, which is the open ball of radius $R$ centered at the origin. We have already shown that this statement is true when the sequence is empty. Assume that the statement is true for the sequence $\{i_{j}\}_{1\leq j\leq m}$. We show that the statement is true for the sequence $\{i_{j}\}_{1\leq j\leq m+1}$.
It suffices to prove that for $x\in B_{R}(0)$,
(32) $\mathcal{D}_{i_{j+1}}\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)=\int_{\mathbb{C}^{N}}\prod_{j=1}^{m+1}a_{i_{j+1}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)$
and that the right hand side is holomorphic over $B_{R}(0)$. Note that $\left\lVert x\right\rVert_{2}<R$, so applying part (e) of Lemma˜2.6 gives that
$\left|\prod_{j=1}^{m+1}a_{i_{j+1}}E_{a}^{\mathcal{R}(\theta)}(x)\right|\leq\sqrt{|H(\mathcal{R})}|\prod_{j=1}^{m+1}a_{i_{j+1}}\exp(\left\lVert a\right\rVert_{2}\left\lVert x\right\rVert_{2}),$
which implies that the right hand side converges. Afterwards, we can use the method used in [matrix]*Proof of Lemma 2.9 to show that the right hand side is holomorphic. To complete the proof, it suffices to show that (32) is true. Without loss of generality, assume that $i_{m+1}=1$.
Suppose $x\in B_{R}(0)$. Observe that
(33) $\partial_{1}\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)=\lim_{\epsilon\rightarrow 0}\int_{\mathbb{C}^{N}}\frac{\prod_{j=1}^{m}a_{i_{j}}(E_{a}(x_{1}+\epsilon,x_{2},\ldots,x_{N})-E_{a}(x))}{\epsilon}d\mu(a).$
By [dunklbound]*Lemma 3.5,
$\frac{|E_{a}(x_{1}+\epsilon,x_{2},\ldots,x_{N})-E_{a}(x)|}{\epsilon}\leq e\sqrt{|H(\mathcal{R})|}\left\lVert a\right\rVert_{2}\exp(\left\lVert a\right\rVert_{2}(\left\lVert x\right\rVert_{2}+\epsilon)).$
If we assume that $\epsilon<R-\left\lVert x\right\rVert_{2}$, then the absolute value of the integrand of the right hand side of (33) is upper bounded by an integrable function, since $\mu$ is exponentially decaying at rate $R$. Then, we can apply the dominated convergence theorem to deduce that
$\partial_{1}\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)=\int_{\mathbb{C}^{N}}\partial_{1}\prod_{j=1}^{m}a_{i_{j}}E_{a}(x)d\mu(a).$
Suppose $\alpha\in\mathcal{R}^{+}$. Then, it is clear that
$\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)-\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(r_{\alpha}x)d\mu(a)$
$=\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}(E_{a}^{\mathcal{R}(\theta)}(x)-E_{a}^{\mathcal{R}(\theta)}(r_{\alpha}x))d\mu(a)$
since $r_{\alpha}x\in B_{R}(0)$. After using the expression given in $\mathcal{D}_{1}$ in Subsection˜2.1, we deduce that
$\mathcal{D}_{1}\int_{\mathbb{C}^{N}}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)=\int_{\mathbb{C}^{N}}\mathcal{D}_{1}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a)=\int_{\mathbb{C}^{N}}a_{1}\prod_{j=1}^{m}a_{i_{j}}E_{a}^{\mathcal{R}(\theta)}(x)d\mu(a),$
which completes the induction.
(E): This is straightforward to deduce from (D). ∎
@@
The papers [matrix, limits_prob_measures] only consider the $A^{N}$ root system. However, (A) and (E) are generalizable to any finite root system after using the method discussed in [limits_prob_measures] that involves applying the results of [dunklbound]. Furthermore, observe that (A) implies (B) and (C). However, we require the exponentially decaying condition to prove (D) and (E).
@@
Define $G_{\mu}^{A^{N-1}(\theta)}$, $c_{\lambda}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\lambda\in\Gamma_{N}$, and $m_{\epsilon}^{A^{N-1}(\theta)}(\mu_{N})\in\mathbb{C}$ for $\epsilon\in\mathcal{P}_{N}$ as they are defined in Lemma˜10.16. Then, the following are equivalent.
@@
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{c_{\lambda}^{A^{N-1}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{c_{\lambda}^{A^{N-1}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  2. (a’)
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m_{\epsilon}^{A^{N-1}(\theta)}(\mu_{N})}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m_{\epsilon}^{A^{N-1}(\theta)}(\mu_{N})}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  3. (a”)
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}G_{\mu_{N}}^{A^{N-1}(\theta)}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}G_{\mu_{N}}^{A^{N-1}(\theta)}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.